# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v2

@@ 1. Introduction
Meaningful social connections play a critical role in young people’s development and well-being (Hawkley and Cacioppo, 2010), and many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]). However, recent trends indicate that social media platforms are increasingly becoming less focused on facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). Initially designed around sharing personal updates with friends, these platforms now often emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s own data illustrates this shift: between 2023 and 2025, the share of time users spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on supporting users “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
@@
This shift away from personal interactions raises concerns about the future of meaningful online relationships, particularly for younger users who have grown up with social media as a central part of their social lives. Most youth today engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite this high engagement, many young users report dissatisfaction with the quality of their social interactions, not because they wish to reduce their usage, but because they expect these platforms to better support their social connections (Kim et al., 2024b). Current platform designs frequently prioritize passive consumption of content over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
@@
Given the central importance of social connections, particularly for youth, and recognizing social media’s vital role in facilitating these interactions, it is increasingly important to explore new design possibilities. Although young people rely heavily on social media, they frequently express dissatisfaction with their current experiences, indicating a significant gap between their social needs and the interactions enabled by existing platforms. Current mainstream social media platforms employ highly standardized design patterns, which limit our collective imagination and contribute to feelings of pessimism among both younger and older users. Consequently, there is a compelling need for innovative designs that complement existing platforms by placing a stronger emphasis on meaningful interpersonal connections. This prompts the critical question of understanding what types of social interactions current designs might be overlooking or inadequately supporting.
@@
To address this question, we conducted a series of co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007). We involved 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants were asked to collaboratively imagine a social media platform specifically designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from the constraints of existing platforms and envision more varied and meaningful social interactions. Additionally, because 3D gaming environments are widely recognized as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated that participants could leverage insights from their gaming experiences to propose novel social media interactions. Specifically, our research aims to answer the following question: What types of interactions do youth familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
@@
Our findings reveal that youth with 3D gaming experiences envision ideal social media platforms not as immersive, navigable environments designed to feel more like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. Together, these visions highlight the youth’s desire for social media that feels more relational, meaningful, and personally rewarding.
@@
We contribute empirical findings on what youth perceive as the most significant shortcomings of current social media platforms, as well as what they desire to experience on social media. Additionally, we propose the concept of spatial integrity for social interactions in online spaces, which encompasses four dimensions of spatial affordances that facilitate meaningful social connections online. These dimensions include spatial presence (the ability to understand the space and identify who is present), spatial composition (the meaning and significance of the space), spatial configuration (the organizational structure of the space), and spatial depth (the extent of information that can be perceived within the space). Furthermore, we demonstrate that the FI method, in the context of social media design, not only prompted creative idea generation and enabled participants to transcend current social media design patterns but also empowered them to envision the positive potential of social media, leaving them feeling more hopeful about its future possibilities.
@@ 2.1.1. Defining Meaningful Social Interaction
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” While these interactions are described as “deeply subjective,” the primary distinction from non-meaningful ones lies in the presence of impact. This impact extends “beyond the immediate interaction itself,” typically involving positive emotions, the exchange of information, or concrete benefits.
The study identified three key types of impact: “emotional impact” (involving feelings, mood changes, or relationship strengthening), “informational impact” (relating to the exchange of knowledge, advice, or understanding), and “tangible impact” (providing concrete benefits such as goods or services). In contrast, non-meaningful interactions were often explicitly described as lacking any such impact.
Factors like the presence of strong ties (e.g., friends and family) and community ties, as well as participation in shared activities, were found to be associated with meaningfulness. However, these appeared to act more as facilitators that enable impact rather than as the primary source of meaning. Other attributes quantitatively linked to greater meaningfulness included whether interactions were planned in advance, memorialized (for example, through photos or videos), and synchronous. Interestingly, the study found that the communication medium itself, whether online or offline, was less important in predicting meaningfulness after accounting for these other characteristics.
While in this paper we treat meaningful social interaction as any interaction that participants say they would actively seek out in an ideal scenario. reflecting their subjective or perceived meaningfulness, we expected these themes to echo in how our participants talked about their experiences.
@@ 2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Social media platforms are often seen as a primary means of connecting with others, especially for young people (Boyd, 2014a). However, despite the widespread use of these platforms, many users report feeling lonely or disconnected (Twenge et al., 2019; Primack et al., 2017). This disconnect highlights a significant gap between the intended purpose of social media and the actual experiences of users.
Social connection is just one of the many reasons people turn to social media, but it remains one of the most important. From staying in touch with close friends to meeting new people or maintaining community ties, users often come to these platforms seeking connection. Mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) support some of these needs, but they are designed to address a wide range of user motivations, including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), even within a single platform, users pursue diverse goals, ranging from building social capital (Phua et al., 2017) to engaging in general media consumption (Menon, 2022). As such, although people may approach these platforms with the intention of connecting socially, they may encounter experiences that prioritize other forms of engagement, potentially leaving their social needs unmet.
This gap highlights the importance of developing a clearer understanding of what it means to support meaningful social connection in online environments. While social media offers both risks, such as social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b), and benefits, such as support for marginalized groups (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and opportunities for identity development and self-expression (Davis, 2012; Lee et al., 2022), social connection remains a central motivation for many users. Particularly for young people, who increasingly rely on digital platforms for social interaction, it is critical to move beyond treating connection as one of many competing platform goals and toward a more focused understanding of how platforms can meaningfully support users’ social relationships.
@@ 2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs that foster meaningful social interactions online, examining both platform-level innovations and individual features. Social media platforms such as BeReal (Kim et al., 2024b), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) have introduced unique affordances for sharing and connecting, with an emphasis on spontaneity, ephemerality, and creative self-expression. This line of research identified different ways to build and maintain relationships through features that prioritize authenticity and playful interaction. However, users have reported increasingly toxic interactions as they gain popularity, with design choices encouraging users to expand their networks and engage in compulsive behaviors, such as tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), which detract from their initial goals of fostering meaningful interactions.
At the feature level, efforts to support meaningful communication include designs that encourage effortful communication (Liu et al., 2021; Zhang et al., 2022) and those that foster feelings of genuine connections (Stepanova et al., 2022). While these features are important, social media experiences are often holistic, involving a complex interplay of multiple design aspects and user behaviors. As a result, such features alone cannot fundamentally alter the overarching narratives or perspectives surrounding social media.
Beyond archetypal social media, recent work has investigated spatial social technologies, which emphasize the value of immersive, embodied interactions. Social virtual reality (VR) platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide users with virtual spaces and embodiment for co-presence and shared experiences, facilitating interactions that feel more personal and tangible compared to 2D environments. Similarly, platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend elements of gaming and social networking, offering users the ability to navigate virtual environments, customize avatars, and participate in dynamic, context-rich group interactions. However, these platforms often are not generally recognized as social media in the generic sense.
What youth envision as ideal social media, both at the broader landscape level and the specific feature level remains unclear, especially when considering the possibilities offered by emerging technologies such as spatial computing platforms. Understanding these perspectives is crucial for reimagining the future of social media and ensuring it meets the evolving needs of its users.
@@ 2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for fostering meaningful social connection, particularly among youth. Large-scale survey research shows that video games serve as a major venue for the development and maintenance of friendships, especially for boys (Lenhart, 2015c). Teens not only play games with friends they know in person but also frequently make new friends through online gameplay (Lenhart, 2015a), with more than half of teens reporting they have met at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay itself, supporting ongoing social relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research also highlights the social and emotional benefits of gaming, including opportunities for cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain game designs promote prosocial outcomes by encouraging players to help and support others, both in-game and beyond (Gentile et al., 2009). Moreover, games are not limited to surface-level entertainment; they can also offer deeper, eudaimonic experiences that involve meaning-making, personal growth, and feelings of connection with others (Daneels et al., 2023). These findings suggest that video games can be powerful tools for fostering meaningful social connections, particularly among youth, and highlight the potential for game-like features to enhance social media experiences.
@@ 2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented challenge in design. While generating a wide range of divergent ideas is a critical first step in innovative design, we often impose unnecessary restrictions on our thinking. This self-limiting cognitive process often confines the design process to address localized problems rather than exploring broader, more optimal solutions.
Social media design is no exception (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Given the ad-based revenue model, platforms are incentivized to increase time spent on their apps (Oth, [n. d.]). However, instead of achieving this through meaningful utility and engagement that genuinely enhances users’ experiences, they often resort to less healthy, compulsive strategies, such as friend recommendations (Kim et al., 2024b) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)). These tactics drive superficial interactions rather than fostering deeper connections, leaving users dissatisfied and reinforcing the perception that all social media designs and user experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
2.3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
FI is an exploratory co-design method that enables participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). At its core, FI creates an environment where people can imagine beyond present limitations by engaging with carefully crafted scenarios. These scenarios typically incorporate immersive elements such as physical artifacts or role-playing activities, with participants taking on generative roles that help them envision futuristic or fictional settings. The method is particularly effective at helping participants overcome design fixation by encouraging them to draw analogies from fictional scenarios and imagine freely outside real-world constraints. While FI shares the exploratory nature of other ideation techniques, it is distinguished by its emphasis on developing futuristic design materials through collaborative dialogue between designers and users (Dindler and Iversen, 2007).
According to Dindler and Iversen (Dindler and Iversen, 2007), implementing this method involves three key steps: first, clearly defining the inquiry’s purpose, whether it’s for staging design situations, exploring future ideas, or driving organizational change; second, developing an appropriate narrative that participants are comfortable with while being distinct enough from current practices to encourage new thinking; and third, creating a compelling plot that introduces tension or conflict within the narrative to motivate specific workshop activities. Unlike traditional role-playing approaches, FI encourages participants to maintain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities while drawing from their personal motivations and capabilities, all while circumventing typical societal constraints.
Further, one of the most powerful ways to overcome design fixation is analogical transfer (Novick, 1988; Buchanan, 1992). We see video games as a rich source of analogies for social media design, offering unique affordances and interaction patterns that can inspire new approaches to fostering meaningful social connections. By examining how video games facilitate social interactions, we can identify design patterns that can be adapted to social media contexts, helping to break free from design fixation and explore novel alternatives.
@@ 3. Method
We conducted co-design interviews with 23 participants between the ages of 15 and 24 to examine how we might envision an ideal social media platform free from real-world constraints and from design fixation on archetypal social media platforms. All participants had prior experience with 3D gaming and familiarity with Harry Potter (Rowling, 2015; har, [n. d.]), a popular series of young adult fantasy novels (and movies based on these stories) in which a subset of the population possesses magical powers. During the interviews, we used the FI method to scaffold participants in generating design ideas.

### v3

@@ 1. Introduction
Meaningful social connections are critical for young people’s development and well-being (Hawkley and Cacioppo, 2010), and social media has great potential to support these. At the individual level, richer and healthier interactions support youth well-being and identity development (Kim et al., 2024c). At the societal level, platforms act as critical infrastructures for sustaining friendships and community ties, making their design choices consequential for social capital (Phua et al., 2017a; Ellison et al., 2007). In fact, many young people turn to social media platforms to build and maintain these relationships (Allen, 2019; Bohn, [n. d.]).
@@
However, social media platforms increasingly shift away from facilitating personal interactions (Nielsen and Fletcher, 2023; Chayka, 2025). While designers initially created these platforms around sharing personal updates with friends, they now emphasize entertainment, news, and engaging larger, less intimate audiences. Meta’s data illustrates this shift: between 2023 and 2025, time spent viewing friends’ content dropped from 22% to 17% on Facebook and from 11% to 7% on Instagram (Bradley et al., 2025). Mark Zuckerberg described this change as aligned with the company’s focus on “the general idea of entertainment and learning about the world and discovering what’s going on” (Chayka, 2025).
@@
This shift raises concerns about meaningful online relationships, particularly for young people who have grown up with social media as central to their social lives. Most young people engage with social media daily, spending substantial time on platforms like Instagram and TikTok (Anderson, 2023). Despite high engagement, many report dissatisfaction with their social interactions, not because they want reduced usage, but because they expect platforms to better support their connections (Kim et al., 2024c). Current designs prioritize passive content consumption over active interaction with close friends (Landesman et al., 2024), contributing to broader concerns about social media’s impact on youth well-being (Oth, 2023).
@@
Given social connections’ importance for young people and social media’s vital role in facilitating them, we must explore new design possibilities. Young people frequently express dissatisfaction with current experiences, revealing a significant gap between their social needs and existing platform interactions. Mainstream platforms employ standardized design patterns that constrain creative possibilities and foster pessimism among users about social media’s potential. This creates a compelling need for innovative designs that complement existing platforms by emphasizing meaningful interpersonal connections. Therefore, we ask the following research question: What types of social interactions might current social media designs be overlooking or inadequately supporting?
@@
To address this question, we conducted co-design interviews using the Fictional Inquiry (FI) method (Dindler and Iversen, 2007) with 23 participants aged 15–24 who have experience with 3D gaming environments like Minecraft (min, [n. d.]) and Roblox (Roblox Corporation, [n. d.]). Participants collaboratively imagined a social media platform designed for students at Hogwarts, the fictional school from Harry Potter (Rowling, 2015; har, [n. d.]). This imaginative context allowed participants to step away from existing platform constraints and envision more varied and meaningful social interactions. Since researchers recognize 3D gaming environments as effective spaces for youth socialization alongside social media (Gentile et al., 2009; Granic et al., 2014; Lenhart, 2015c), we anticipated participants could leverage gaming insights to propose novel social media interactions. Specifically, our research addresses: What types of interactions do young people familiar with 3D gaming environments envision as ideal for supporting meaningful remote social connections?
@@
Our findings reveal that young people with 3D gaming experience envision ideal social media platforms as immersive, navigable environments that designers should create to feel like neighborhoods, community centers, and shared social spaces. Participants described six priorities: (1) intuitive social navigation modeled on real-world movement and presence; (2) shared experiences through collaborative activities and casual co-presence; (3) communal environments that prioritize close relationships, conversations, and low-pressure interaction; (4) flexible self-presentation through avatars, personal spaces, and contextual identity management; (5) intentional engagement that replaces passive scrolling with purposeful navigation; and (6) playful, low-stakes social mechanics inspired by gaming. These visions highlight young people’s desire for social media that feels more relational, meaningful, and personally rewarding.
@@
We contribute empirical findings on young people’s perceptions of current social media platforms’ shortcomings and their desired experiences. We propose the concept of spatial integrity for social interactions in online spaces, encompassing four dimensions of spatial affordances that facilitate meaningful social connections: spatial presence (understanding the space and identifying who is present), spatial composition (the space’s meaning and significance), spatial configuration (the space’s organizational structure), and spatial depth (the extent of perceivable information within the space). We demonstrate that the FI method in social media design contexts prompts creative idea generation, enables participants to transcend current design patterns, and empowers them to envision social media’s positive potential, leaving them more hopeful about future possibilities.
@@ 2.1.1. Defining Meaningful Social Interaction
Litt et al. (Litt et al., 2020) define meaningful social interactions as “interactions that people believe enhance their lives, the lives of their interaction partners, or their relationships, with emotional, informational, or tangible impact.” Though “deeply subjective,” these interactions are distinguished by their impact “beyond the immediate interaction itself,” involving positive emotions, information exchange, or concrete benefits.
The study identified three key impact types: “emotional impact” (feelings, mood changes, or relationship strengthening), “informational impact” (knowledge, advice, or understanding exchange), and “tangible impact” (concrete benefits like goods or services). Non-meaningful interactions lacked any such impact.
Strong ties (friends and family), community ties, and shared activities were associated with meaningfulness, but acted as facilitators enabling impact rather than primary sources of meaning. Other attributes linked to meaningfulness included advance planning, memorialization (photos or videos), and synchronicity. The communication medium itself—online or offline—was less important than these other characteristics.
We define meaningful social interaction as any interaction that participants would actively seek in an ideal scenario, reflecting their subjective meaningfulness. We expected these themes to emerge in participants’ descriptions of their experiences.
@@ 2.1.2. Social Connection as a Key User Need Left Unmet by Today’s Platforms
Though social media platforms are primary connection tools for young people (Boyd, 2014a), many users report feeling lonely or disconnected despite widespread use (Twenge et al., 2019; Primack et al., 2017), highlighting a gap between intended purpose and actual experience.
While social connection remains a primary motivation for social media use, mainstream platforms like Instagram (All, [n. d.]), YouTube (you, [n. d.]), and Discord (discord page, [n. d.]) are designed for diverse user motivations including entertainment, information seeking, and self-promotion (Whiting and Williams, 2013; Menon, 2022). From a uses and gratifications perspective (Katz et al., 1973), users pursue varied goals from building social capital (Phua et al., 2017b) to media consumption (Menon, 2022). Users seeking social connection may encounter experiences prioritizing other engagement forms, leaving social needs unmet.
This gap underscores the need for a clearer understanding of meaningful social connections in online environments. While social media presents risks like social comparison and privacy concerns (Yau and Reich, 2019; Weinstein and James, 2022; Kim et al., 2024a; Boyd, 2014b) alongside benefits like marginalized group support (Li et al., 2023; Bates et al., 2020; Steeds et al., 2025) and identity development opportunities (Davis, 2012; Lee et al., 2022), social connection remains central for many users. For young people increasingly relying on digital platforms, we must move beyond treating connection as one competing goal toward a focused understanding of how platforms can meaningfully support social relationships.
@@ 2.2.1. Designing for Meaningful Social Connections
HCI research has extensively explored designs fostering meaningful social interactions online at both platform and feature levels. Platforms like BeReal (Kim et al., 2024c), Snapchat (Bayer et al., 2016), Miitomo (Kasunic and Kaufman, 2017), and TikTok (Barta and Andalibi, 2021; Schaadhardt et al., 2023) introduced unique affordances emphasizing spontaneity, ephemerality, and creative self-expression. This research identified ways to build relationships through features prioritizing authenticity and playful interaction. However, users report increasingly toxic interactions as platforms gain popularity, with design choices encouraging network expansion and compulsive behaviors like tracking Snap scores (Chambers and King, 2022; van Essen and Van Ouytsel, 2023), detracting from meaningful interaction goals.
Feature-level efforts include designs encouraging effortful communication (Liu et al., 2021; Zhang et al., 2022) and genuine connections (Stepanova et al., 2022). However, social media experiences are holistic, involving a complex interplay of multiple design aspects and user behaviors. Individual features alone cannot fundamentally alter overarching social media narratives.
Beyond archetypal social media, recent work investigates spatial social technologies emphasizing immersive, embodied interactions. Social VR platforms (Zamanifard and Freeman, 2023; Freeman et al., 2024, 2020; Zamanifard and Freeman, 2019) provide virtual spaces and embodiment for co-presence and shared experiences, facilitating more personal and tangible interactions than 2D environments. Platforms like Gather.town (Duarte et al., 2023; Tu, 2022) and VRChat (Chen et al., 2024; Deighan et al., 2023; Rzeszewski and Evans, 2024) blend gaming and social networking, enabling virtual environment navigation, avatar customization, and dynamic group interactions. However, these platforms are not generally recognized as conventional social media.
What youth envision as ideal social media at the landscape and feature levels remains unclear, especially considering emerging technologies like spatial computing platforms. Understanding these perspectives is crucial for reimagining social media’s future to meet users’ evolving needs.
@@ 2.2.2. Video Games as a Source of Meaningful Social Connection
Video games are increasingly recognized as valuable environments for meaningful social connection, particularly among youth. Large-scale surveys show games serve as major venues for friendship development and maintenance, especially for boys (Lenhart, 2015c). Teens play with known friends and make new ones through online gameplay (Lenhart, 2015a), with over half reporting meeting at least one friend online (Lenhart, 2015a, b). These interactions extend beyond gameplay, supporting ongoing relationships through shared experiences, communication, and collaboration (Lenhart, 2015c).
Research highlights gaming’s social and emotional benefits, including cooperation, empathy, and prosocial behavior (Gentile et al., 2009; Granic et al., 2014). Certain designs promote prosocial outcomes by encouraging players to help others in-game and beyond (Gentile et al., 2009). Games offer deeper, eudaimonic experiences involving meaning-making, personal growth, and connection (Daneels et al., 2023). These findings suggest video games can foster meaningful social connections among youth and highlight the potential for game-like features in social media.
@@ 2.3. Recognizing and Overcoming Social Media Design Fixation
Design fixation (Jansson and Smith, 1991)—the tendency to remain constrained by conventional thinking or existing solutions—is a well-documented design challenge. Though generating divergent ideas is critical for innovation, we often impose unnecessary restrictions. This self-limiting process confines design to localized problems rather than exploring broader, optimal solutions.
Social media design faces similar fixation (Sundaram, [n. d.]; Pardes, 2020; Pitt, 2021; Nielsen and Fletcher, 2023; VanDyke, 2020). Ad-based revenue models incentivize increased time on apps (Oth, [n. d.]). Platforms use compulsive strategies like friend recommendations (Kim et al., 2024c) or engagement metrics (e.g., Snap scores (Rozgonjuk et al., 2021)), creating distrustful user communities (Kim et al., 2025). Rather than proactively exploring new paradigms of designs and interactions, much recent work on social media design has disproportionately focused on reactively addressing existing harm (Kim et al., 2024b). The domination of mainstream platforms along with this reactive approach to design leaves users dissatisfied and reinforces perceptions that all social media experiences are fundamentally the same (Sundaram, [n. d.]; Pardes, 2020).
Analogical transfer offers one powerful method for overcoming design fixation (Novick, 1988; Buchanan, 1992). Therefore, in this paper, we attempt to break out of the design and cognitive fixation that assumes the future of social media is bleak by freeing participants from real-world constraints and beliefs and leveraging elements from fictional settings to serve as sources for analogical transfer.
@@ 3. Method
We conducted co-design interviews with 23 participants aged 15-24 to examine how we might envision ideal social media platforms free from real-world constraints and design fixation on archetypal platforms. All participants had 3D gaming experience and Harry Potter (Rowling, 2015; har, [n. d.]) familiarity, a popular young adult fantasy series where some possess magical powers. We used the FI method to scaffold participants’ design idea generation. All procedures were approved by the university’s Institutional Review Board (IRB).
3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
FI is an exploratory co-design method enabling participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). FI creates environments where people imagine beyond present limitations through carefully crafted scenarios incorporating immersive elements like physical artifacts or role-playing activities. Participants take generative roles envisioning futuristic or fictional settings. The method effectively overcomes design fixation by encouraging analogies from fictional scenarios and free imagination outside real-world constraints. Unlike other ideation techniques, FI emphasizes developing futuristic design materials through collaborative designer-user dialogue (Dindler and Iversen, 2007).
Dindler and Iversen (Dindler and Iversen, 2007) outline three key implementation steps: defining the inquiry’s purpose (staging design situations, exploring future ideas, or driving organizational change); developing an appropriate narrative that participants find comfortable yet distinct enough from current practices to encourage new thinking; and creating compelling plots introducing tension or conflict to motivate workshop activities. Unlike traditional role-playing, FI encourages participants to maintain their identities and expertise while operating in fictional contexts, exploring new possibilities through personal motivations and capabilities while circumventing societal constraints.
We view video games as rich analogy sources for social media design, offering unique affordances and interaction patterns inspiring new approaches to meaningful social connections. By examining how games facilitate social interactions, we can identify adaptable design patterns for social media contexts, breaking free from fixation to explore novel alternatives.